what is the way and answer

(2X-5) TO THE FOURTH POWER

Answer by sdmmadam@yahoo.com(530)

what is the way and answer
(2X-5) TO THE FOURTH POWER
Required to find the fourth power of (2x-5)
We use the simple concept: (a)^4 = (a^2)^2
So, we write
(2x-5)^4
= [(2x-5)^2]^2
= [(2x)^2+(-5)^2 + 2X(2x)X(-5)]^2 (using (a+b)^2 = a^2 +b^2 +2ab)
=(4x^2+25-20x)^2
=(4x^2)^2+(25)^2+(-20x)^2+2X(4x^2)X(25)+2X(25)X(-20x)+2X(-20x)X(4x^2)
[using (a+b+c)^2 = a^2 + b^2 + c^2 +2ab+2bc+2ca
and here a=4x^2,b=25 and c=(-20x)]
=16x^4+625+400x^2+200x^2-1000x-160x^3
=16x^4-160x^3+600x^2-1000x+625